OpenStudy (anonymous):

what is the length of the segment connecting points (1,2) and (5,5) in the xy-coordinate plane

OpenStudy (anonymous):

5

OpenStudy (anonymous):

it is a right triangle, and its a 3-4-5 triangle at that. The five is the hypotenuse

OpenStudy (anonymous):

\[d(P_{1},P_{2})=\sqrt{(x _{1}-x _{2})^{2}+(y _{1}-y _{2})^{2}}\] \[d((1,2),(5,5))=\sqrt{(1 -5 )^{2}+(2 -5)^{2}}\] \[d((1,2),(5,5))=\sqrt{(-4)^{2}+(-3)^{2}}\] \[d((1,2),(5,5))=\sqrt{16+9}\] \[d((1,2),(5,5))=\sqrt{25}\] \[d((1,2),(5,5))=5\]